OpenStudy (anonymous):

Will award a medal, (dy/dx)+(y/3)=0 find the equation where when y=10 x=0. I had this problem for a calc II quiz today and got y=10e^(-x^2/6)

myininaya (myininaya):

looks like a seperation of variables question \[\frac{dy}{dx}=\frac{-y}{3} \\ \frac{dy}{y}=\frac{-1}{3} dx\] integrate both sides

OpenStudy (anonymous):

yea so you get ln|y|=-x^2/6+c

myininaya (myininaya):

not exactly

myininaya (myininaya):

\[\ln|y|=\frac{-1}{3}x+C\]

myininaya (myininaya):

you didn't have an x next to -1/3

myininaya (myininaya):

you know before integrating

OpenStudy (anonymous):

bummer hopefully i get some partial credit

myininaya (myininaya):

if you had \[\frac{dy}{y}=\frac{-1}{3}x dx \text{ then you would have } \ln|y|=\frac{-1}{3} \frac{x^2}{2} +C\] which is what you got

myininaya (myininaya):

it is possible you will get possible credit

myininaya (myininaya):

partial*

myininaya (myininaya):

you did integrate the other side correctly :)

OpenStudy (anonymous):

hahaha